# On the variational dual formulation of the Nash system and an adaptive convex gradient-flow approach to nonlinear PDEs

### v1

@@ 1. Introduction
Here the unknown function is $\psi=\psi(\tau,x)$, where $\tau\in[0,T]$ and $x$ belongs to the periodic box $\mathbb{T}^{N}$. The system is complemented with the terminal condition
@@
In this paper, we are concerned with the variational dual formulation of such nonlinear systems, with an emphasis on the noise-free Nash system. The idea originates with Brenier [14, 15], who proposed, for the incompressible Euler equations and related models, to study solutions that minimize the Lagrangian action, i.e., the time integral of the kinetic energy (for recent applications of the least action principle as a selection criterion for systems of hyperbolic conservation laws, see [18, 19]). Although this minimization problem may fail to have a solution, it leads to a dual problem with more favorable convexity properties. Brenier derived an explicit relation linking smooth solutions of the incompressible Euler equations on short time intervals with the variational dual solutions. In [16], this technique was applied to the multi-stream Euler-Poisson system. A numerical implementation for the dual variational formulation of the quadratic porous medium equation and the Burgers equation has recently been done in [25]. In [30, 31], the first author extended Brenier’s approach by identifying structures in nonlinear PDEs that allow variational dual formulation with favorable properties. These dual problems are related to optimal transport, specifically its ballistic variant [31, 25].
@@
The primary goal of these approaches is to define convex variational principles for conservative and non-conservative (including dissipative) physical models to facilitate their solution, whether exact or approximate (through computation). For instance, they allow, using methods of the calculus of variations, the definition and proofs of existence of variational dual solutions to the Euler-Lagrange PDE systems of physical energy functionals whose global minimizers do not exist [28, Sec. 5], [6]. However, a somewhat unsettling aspect of the duality scheme with the zero base state is that many consistency results are obtained only on small time intervals, cf. [14, Theorems 2.3 and 3.1], [30, Theorem 3.8], [16, Theorem 5.1]. In [31], the first author succeeded in extending the consistency to large time intervals by introducing suitable time-dependent weights; as an application, he derived a variant of Dafermos principle for various PDE models that admit the variational dual formulation.
@@
Subsequently, in Section 3 we propose a scheme, based on Hilbertian gradient flows and suitable for a wide class of nonlinear PDEs, that produces, from an initial base state, a sequence of new base states anticipated to converge to a solution of the original PDE. The proposed adaptive gradient flow scheme is a natural outgrowth of the practical experience with computational schemes based on Newton’s method for approximating the Euler-Lagrange system of the dual variational principle. From the very first nonlinear problem solved in [22] (Euler’s ODE system for dynamics of a rigid body), it was clear that the idea of base states was crucial to the practical success of the scheme, further confirmed in [23] in the context of (inviscid) Burgers (see also [25]) and in [28] for nonconvex elastostatics and dynamics (in space dimension $1$). In [24], base state changes with step-size control in a Newton approximation scheme were used, and the possibility of a gradient flow-based approach was alluded to. The gradient flow scheme formalized here in detail is the logical natural development of those ideas in an exact setting, i.e., without approximation.
@@ 2.1. The variational dual formulation of the Nash system
where the infimum is taken over all pairs $(z,M)\in L^{2}((0,T)\times\Omega;\mathbb{R}^{n})\times L^{1}((0,T)\times\Omega;\mathbb{R}^{n\times n}_{s})$. Note that $\mathcal{K}(Q,B)$ is a negative-semidefinite quadratic form w.r.t. $Q$ for fixed $B$. Mimicking [30, Remark 3.6], it is possible to see that (2.11) is equivalent to
@@ 2.2. Consistency and existence of solutions
Although we put our main attention to the Nash system, the results below can be easily extended to other quadratic PDEs, for instance, the ones studied in [14, 30].
@@
In the pioneering work [14], Brenier considered the variational dual formulation (with zero base state) for the quadratic Burgers equation — or, equivalently, to the one-dimensional quadratic H–J — which corresponds to $n=N=p=1$ and $\bar{v}=0$ in our framework. He shows explicitly that, in that setting, the dual solution $(\rho v,\rho)$ is merely a vector-valued Radon measure (see also Remark 2.12 below, which explains the loss of regularity) and that the support of the measure $\rho$ is strictly smaller than $0,T]\times\Omega$; here $\rho$ can be interpreted as the probability measure transported by $v$ whose marginal at time $T$ is the Lebesgue measure on $\Omega$ (cf. the proof of Corollary 2.6). Consequently, $v$ cannot be recovered as the Radon-Nikodym derivative of $\rho v$ w.r.t. $\rho$ outside of the support of $\rho$. Corollary 2.6 shows that a small perturbation of the base state $\bar{v}=0$ can dramatically improve the regularity of the dual solution and secure one-to-one correspondence between $v$ and the variational dual solution. The same observations apply in the case $N=1$ with $p>1$; under these conditions, the existence of a variational dual solution to H-J with $\bar{v}=0$ — again a vector-valued Radon measure — follows from [30, Theorem 4.6].
@@
$(\mathfrak{U}^{*}(y))_{ijl,pqr}=0\ \mathrm{for}\ \mathrm{all}\ \mathrm{other}\ \mathrm{entries}.$
@@
$\operatorname{Tr}\mathfrak{U}^{*}(y)=\frac{d}{2}\sum_{i=1}^{N}y_{i}.$ (2.29)
@@
$J(v_{0},\bar{v},T)\geq\int_{0}^{T}\left[-(v_{0},\Pi\bar{v})+\frac{1}{2}(\bar{v},\bar{v})\right]\,dt+\mathcal{K}(\bar{v}-\Pi\bar{v},0)\\ =\int_{0}^{T}\left[-(\Pi v_{0},\Pi\bar{v})+\frac{1}{2}(\bar{v},\bar{v})\right]\,dt-\frac{1}{2}\int_{0}^{T}(\bar{v}-\Pi\bar{v},\bar{v}-\Pi\bar{v})\,dt\\ =\int_{0}^{T}\left[-(\Pi v_{0},\Pi\bar{v})+\frac{1}{2}(\Pi\bar{v},\Pi\bar{v})\right]\,dt=\mathfrak{C}(\Pi\bar{v},\Pi v_{0}).$
@@ 3. The utility of base states in practical considerations: A formal convex gradient flow scheme
In this section we describe a formal scheme for obtaining weak solutions to a class of PDE; the class includes many equations of continuum mechanics. The scheme is quite generally applicable, up to minor adjustments, including the case of the Nash system, see Section 5 below. Comparable adaptive numerical schemes employing the idea of variational dual solutions and adjustable base states for solving nonlinear differential equations have recently been demonstrated in [22, 28, 23, 24]. These methods were developed in connection with the nonlinear system of ODE of Euler for the dynamics of a rigid body, nonconvex elastostatic and dynamics of a bar (without regularization), (inviscid) Burgers equation, and the problem of traveling waves of a dispersive, nonlocal, nonlinear semi-discrete Burgers equation.
@@ 3.1. Variational setup
where $\Omega$ is a given bounded domain in $\mathbb{R}^{m}$ with sufficiently regular boundary $\partial\Omega\supset\bigcup_{\Gamma}\partial\Omega_{\Gamma}$, upper-case Greek indices index the number of equations. Here, ${\mathcal{C}}\in\mathbb{R}^{d\times u}$, ${\mathcal{F}},G,V$ are given functions of their argument, and $U^{(0)},B$ are specified functions.
@@
The $(t,x)$ dependence of $H$, introduced in [1, Sec. 5] with the stated goal of aiding in the recovery of non-unique solutions of the primal problem, includes the deployment of “base states” in the design of the dual functional [^2] As an example, in [23, Sec. 5.4] piecewise-in-time solutions of Burgers with small viscosity are used as base states to recover entropy solutions to inviscid Burgers, when written in H-J form.. These are prescribed fields $(t,x)\mapsto\bar{U}(t,x)$. We now require that the choice of $H$, given a base state $\bar{U}$, be such that there exists a non-empty, open, convex subset $\mathcal{O}\subset\mathbb{R}^{l}$ and a function $U^{(H)}:\mathcal{O}\times\mathbb{R}^{u}\to\mathbb{R}^{u}$ satisfying
@@
    satisfies $\mathcal{D}(t,x)\subset\mathcal{O}$,
@@
Before proceeding further, we first show that choices of $H(U,\bar{U})$ satisfying (3.4) are easily made [^3] From now on, we employ a small abuse of notation and write $H(U,\bar{U})$, $\mathcal{L}_{H}(U,\mathcal{D},\bar{U})$ instead of $H(U,t,x)$, $\mathcal{L}_{H}(U,\mathcal{D},t,x)$, etc.; for instance, suppose
@@
[…]
Since the set of positive-definite matrices is an open subset of $\mathbb{R}^{u\times u}$, assuming $\mathcal{L}_{H}$ is $C^{2}$ in a neighborhood of $(\bar{U},0;\bar{U})\in\mathbb{R}^{u}\times\mathbb{R}^{l}$ (for the last argument fixed), by the Implicit function theorem, there exists a (w.l.o.g. convex) neighborhood $\mathcal{O}\subset\mathbb{R}^{l}$ containing $\mathcal{D}=0$ and a mapping $U^{(H)}(\cdot\,,\bar{U}):\mathcal{O}\to\mathbb{R}^{u}$ satisfying (3.4a), for each fixed choice of $\bar{U}\in\mathbb{R}^{u}$.
@@
(noting that $\mathcal{L}_{H}$ depends on the field $U$ only through its pointwise values). Due to the affineness of $\widehat{S}$ in $D$, the functional $\widetilde{S}[\ \cdot\ ;\bar{U}]$ is weakly lower-semicontinuous and convex in the Hilbert space
@@
(Rigorously speaking, $\widehat{S}_{H}[U,D]$ has a certain domain in $\mathcal{H}$ where $\mathcal{D}$ and the boundary traces are well-defined, and the Dirichlet boundary condition (3.5) holds; outside this domain, we simply set $\widehat{S}_{H}[U,D]:=+\infty$.) Furthermore, for a.e. $(t,x)\in[0,T]\times\Omega$, there exists a $\mathbb{R}^{l}\times\mathbb{R}^{u}$ - neighborhood centered in $(0,\bar{U}(t,x))$ in which $\mathcal{L}_{H}\big(\cdot,\mathcal{D}(t,x),\bar{U}(t,x)\big)$ is strictly concave in $U$ w.r.t. its first argument. Consequently, for a given base state $\bar{U}:[0,T]\times\Omega\to\mathbb{R}^{u}$, there exists a convex set $\mathcal{O}^{*}_{\bar{U}}$ in $\mathcal{H}$ containing zero such that for every $D\in\mathcal{O}^{*}_{\bar{U}}$ one has
@@
In other words, for each fixed base state $\bar{U}:[0,T]\times\Omega\to\mathbb{R}^{u}$ there exists a convex set $\mathcal{O}^{*}_{\bar{U}}$ in $\mathcal{H}$ in which the dual functional $S_{H}[D;\bar{U}]$ is convex and its Euler-Lagrange equations and side conditions are given by (3.1), with the substitution $U(t,x)\to U^{(H)}(\mathcal{D}(t,x),\bar{U}(t,x))$. The DtP map $(\mathcal{D},\bar{U})\mapsto U^{(H)}(\mathcal{D},\bar{U})$ is well-defined for $D\in\mathcal{O}^{*}_{\bar{U}}$, with the maximizer in (3.8) being an isolated critical point of $\mathcal{L}_{H}(\,\cdot\,,\mathcal{D}(t,x);\bar{U}(t,x))$.
@@
[…]
    so that
@@
    or the flow reaches a fake time $\tilde{\mathsf{s}}^{*}_{1}$ when (3.14) is violated.
@@
    In cases where the solutions to (3.1) are not unique, we can find only one of them. However, by varying the initial base state $\bar{U}^{1}$, it is expected to be possible to capture additional solutions as well. In principle, any solution can be retrieved, using observation (3) at the end of Sec. 3.1. In practice, see the approximation of a) a wide variety of traveling waves of a dispersive semi-discrete Burgers equation in [24] (with an algorithm which does not exploit the feature (3.13)), and b) even unstable solutions of the primal problem in a stable manner in [28, Sec. 6.3.4] that deals with the nonconvex elastodynamics of a bar (utilizing only one base state and stage).
@@
It is natural to expect that a starting base state in time-dependent problems can be more effective if the real-time interval $[0,T]$ is divided into smaller disjoint sub-intervals whose closed union covers $[0,T]$, with the dual problem solved by the above gradient flow (in fake time) in each such sub-interval, defined with the real-time initial condition generated by the solution of the previous sub-interval. Thus, in this strategy, base states are switched both in real-time (to provide constant-in-real-time starting base state guesses for each sub-interval, at a minimum), as well as potentially in fake-time for the gradient flow within each sub-interval of real time. In experience with nonlinear transient problems computed without the gradient flow strategy [22, Sec. 5.3]-[23], such switching of base states across sub-intervals in real-time was found to be essential for success (as well as computational efficiency). On the other hand, [24] involves a time-independent problem where switching base states between stages of Newton method based iterations (the analog of the gradient flow stages) was found to be essential.

### v2

@@ Abstract.
Keywords: convex duality, generalized optimal transport, gradient flow
MSC [2020] 35D99, 35L40, 37K58, 41A60, 49Q99
@@ 1. Introduction
Here [^1] The Einstein summation convention is not used here but will be adopted in Section 3. the unknown function is $\psi=\psi(\tau,x)$, where $\tau\in[0,T]$ and $x$ belongs to the periodic box $\mathbb{T}^{N}$. The system is complemented with the terminal condition
@@
In this paper, we are concerned with the variational dual formulation of such nonlinear systems, with an emphasis on the noise-free Nash system. The idea originates with Brenier [14, 15], who proposed, for the incompressible Euler equations and related models, to study solutions that minimize the Lagrangian action, i.e., the time integral of the kinetic energy (for recent applications of the least action principle as a selection criterion for systems of hyperbolic conservation laws, see [19, 20]). Although this minimization problem may fail to have a solution, it leads to a dual problem with more favorable convexity properties. Brenier derived an explicit relation linking smooth solutions of the incompressible Euler equations on short time intervals with the variational dual solutions. In [16], this technique was applied to the multi-stream Euler-Poisson system. A numerical implementation for the dual variational formulation of the quadratic porous medium equation and the Burgers equation has recently been done in [26]. In [31, 32], the first author extended Brenier’s approach by identifying structures in nonlinear PDEs that allow variational dual formulations with favorable properties. These dual problems are related to optimal transport, specifically its ballistic variant [32, 26]. Notably, the functionals arising in this context are not expected to be self-dual, so the formalism of [18] is not applicable.
@@
The primary goal of these approaches is to define convex variational principles for conservative and non-conservative (including dissipative) physical models to facilitate their solution, whether exact or approximate (through computation). For instance, they allow, using methods of the calculus of variations, the definition and proofs of existence of variational dual solutions to the Euler-Lagrange PDE systems of physical energy functionals whose global minimizers do not exist [29, Sec. 5], [6]. However, a somewhat unsettling aspect of the duality scheme with the zero base state is that many consistency results are obtained only on small time intervals, cf. [14, Theorems 2.3 and 3.1], [31, Theorem 3.8], [16, Theorem 5.1]. In [32], the first author succeeded in extending the consistency to large time intervals by introducing suitable time-dependent weights; as an application, he derived a variant of Dafermos principle for various PDE models that admit the variational dual formulation.
@@
Subsequently, in Section 3 we propose a novel scheme, based on Hilbertian gradient flows and suitable for a wide class of nonlinear PDEs, that produces, from an initial base state, a sequence of new base states anticipated to converge to a solution of the original PDE. The proposed adaptive gradient flow scheme is a natural outgrowth of the practical experience with computational schemes based on Newton’s method for approximating the Euler-Lagrange system of the dual variational principle. From the very first nonlinear problem solved in [23] (Euler’s ODE system for dynamics of a rigid body), it was clear that the idea of base states was crucial to the practical success of the scheme, further confirmed in [24] in the context of (inviscid) Burgers (see also [26]) and in [29] for nonconvex elastostatics and dynamics (in space dimension $1$). In [25], base state changes with step-size control in a Newton approximation scheme were used, and the possibility of a gradient flow-based approach was alluded to. The gradient flow scheme formalized here in detail constitutes a new and logically natural development of these ideas. We work in an exact setting, i.e., without approximation.
@@ 2.1. The variational dual formulation of the Nash system
where the infimum is taken over all pairs $(z,M)\in L^{2}((0,T)\times\Omega;\mathbb{R}^{n})\times L^{1}((0,T)\times\Omega;\mathbb{R}^{n\times n}_{s})$. Note that $\mathcal{K}(Q,B)$ is a negative-semidefinite quadratic form w.r.t. $Q$ for fixed $B$. Mimicking [31, Remark 3.6], it is possible to see that (2.11) is equivalent to
@@ 2.2. Consistency and existence of solutions
Although our main focus is on the noise-free Nash system, we are confident that the results below can be extended to other quadratic PDEs, for instance, to those studied in [14, 31].
@@
In the pioneering work [14], Brenier considered the variational dual formulation (with zero base state) for the quadratic Burgers equation — or, equivalently, to the one-dimensional quadratic H–J — which corresponds to $n=N=p=1$ and $\bar{v}=0$ in our framework. He shows explicitly that, in that setting, the dual solution $(\rho v,\rho)$ is merely a vector-valued Radon measure (see also Remark 2.12 below, which explains the loss of regularity) and that the support of the measure $\rho$ is strictly smaller than $0,T]\times\Omega$; here $\rho$ can be interpreted as the probability measure transported by $v$ whose marginal at time $T$ is the Lebesgue measure on $\Omega$ (cf. the proof of Corollary 2.6). Consequently, $v$ cannot be recovered as the Radon-Nikodym derivative of $\rho v$ w.r.t. $\rho$ outside of the support of $\rho$. Corollary 2.6 shows that a small perturbation of the base state $\bar{v}=0$ can dramatically improve the regularity of the dual solution and secure one-to-one correspondence between $v$ and the variational dual solution. The same observations apply in the case $N=1$ with $p>1$; under these conditions, the existence of a variational dual solution to H-J with $\bar{v}=0$ — again a vector-valued Radon measure — follows from [31, Theorem 4.6].
@@
$(\mathfrak{U}^{*}(y))_{ijl,hqr}=0\ \mathrm{for}\ \mathrm{all}\ \mathrm{other}\ \mathrm{entries}.$
@@
$\operatorname{Tr}\mathfrak{U}^{*}(y)=\frac{p}{2}\sum_{i=1}^{N}y_{i}.$ (2.29)
@@
$\mathcal{J}(v_{0},\bar{v},T)\geq\int_{0}^{T}\left[-(v_{0},\Pi\bar{v})+\frac{1}{2}(\bar{v},\bar{v})\right]\,dt+\mathcal{K}(\Pi\bar{v}-\bar{v},0)\\ =\int_{0}^{T}\left[-(\Pi v_{0},\Pi\bar{v})+\frac{1}{2}(\bar{v},\bar{v})\right]\,dt-\frac{1}{2}\int_{0}^{T}(\bar{v}-\Pi\bar{v},\bar{v}-\Pi\bar{v})\,dt\\ =\int_{0}^{T}\left[-(\Pi v_{0},\Pi\bar{v})+\frac{1}{2}(\Pi\bar{v},\Pi\bar{v})\right]\,dt=\mathfrak{C}(\Pi\bar{v},\Pi v_{0}).$
@@ 3. The utility of base states in practical considerations: A formal convex gradient flow scheme
In this section we describe a formal scheme for obtaining weak solutions to a class of PDEs; the class includes many equations of continuum mechanics. The scheme is quite generally applicable, up to minor adjustments, including the case of the Nash system, see Section 5 below. Comparable adaptive numerical schemes employing variational dual solutions and adjustable base states — but not based on the gradient flow idea developed here — for solving nonlinear differential equations have recently been demonstrated in [23, 29, 24, 25]. These methods were developed in connection with the nonlinear system of ODEs of Euler for the dynamics of a rigid body, nonconvex elastostatic and dynamics of a bar (without regularization), (inviscid) Burgers equation, and the problem of traveling waves of a dispersive, nonlocal, nonlinear semi-discrete Burgers equation.
@@ 3.1. Variational setup
where $\Omega$ is a given bounded domain in $\mathbb{R}^{m}$ with sufficiently regular boundary $\partial\Omega\supset\bigcup_{\Gamma}\partial\Omega_{\Gamma}$ and unit normal vector $n$, and the upper-case Greek indices index the number of equations. Here, ${\mathcal{C}}\in\mathbb{R}^{d\times m}$ is a fixed matrix, and ${\mathcal{F}},G,V,B$ and $U^{(0)}$ are, respectively, $\mathbb{R}^{d\times m}$-, $d$-,$d$-, $\mathbb{R}^{d\times m}$- and $u$-dimensional given functions of their argument.
@@
The $(t,x)$ dependence of $H$, introduced in [1, Sec. 5] with the stated goal of aiding in the recovery of non-unique solutions of the primal problem, includes the deployment of “base states” in the design of the dual functional [^3] As an example, in [24, Sec. 5.4] piecewise-in-time solutions of Burgers with small viscosity are used as base states to recover entropy solutions to inviscid Burgers, when written in H-J form.. These are prescribed fields $(t,x)\mapsto\bar{U}(t,x)$. We now require that the choice of $H$, given $t,x$ and a base state $\bar{U}$, be such that there exists a non-empty, open, convex subset $\mathcal{O}=\mathcal{O}(t,x)\subset\mathbb{R}^{l}$ and a function $U^{(H)}:\mathcal{O}\times\mathbb{R}^{u}\to\mathbb{R}^{u}$ satisfying
@@
    satisfies $\mathcal{D}(t,x)\subset\mathcal{O}(t,x)$,
@@
Before proceeding further, we first show that choices of $H(U,\bar{U})$ satisfying (3.4) are easily made [^4] From now on, we employ an abuse of notation and write $H(U,\bar{U})$, $\mathcal{L}_{H}(U,\mathcal{D},\bar{U})$ instead of $H(U,t,x)$, $\mathcal{L}_{H}(U,\mathcal{D},t,x)$, etc.; for instance, suppose
@@
$\partial_{UU}\mathcal{L}_{H}(\bar{U},0,\bar{U})>0$ (3.8)
in the sense of symmetric, positive-definite matrices.
By the implicit function theorem [^5] Remember that here we work at a formal level and ignore regularity issues., there exists a (w.l.o.g. convex) neighborhood $\mathcal{O}=\mathcal{O}(t,x)\subset\mathbb{R}^{l}$ containing $\mathcal{D}=0$ and a mapping $U^{(H)}(\cdot\,,\bar{U}):\mathcal{O}\to\mathbb{R}^{u}$ satisfying (3.4a), for each fixed choice of $t,x$ and $\bar{U}\in\mathbb{R}^{u}$.
@@
(noting that $\mathcal{L}_{H}$ depends on the field $U$ only through its pointwise values). Since $\widehat{S}$ is affine [^6] Rigorously speaking, $\widehat{S}_{H}[U,\cdot]$ has a certain convex domain in $\mathcal{H}$ where $\mathcal{D}$ and the boundary traces are well-defined, and the Dirichlet boundary condition (3.5) holds; outside this domain, we simply set $\widehat{S}_{H}[U,D]:=+\infty$; this is enough to guarantee the convexity of $\widetilde{S}_{H}$. in $D$, the functional $\widetilde{S}_{H}[\ \cdot\ ;\bar{U}]$ is convex in the Hilbert space
@@
Since the set of positive-definite matrices is open in $\mathbb{R}^{u\times u}$, using (3.8) we infer that for a.e. $(t,x)\in[0,T]\times\Omega$ there exists a $\mathbb{R}^{l}\times\mathbb{R}^{u}$-neighborhood centered in $(0,\bar{U}(t,x))$ in which $\mathcal{L}_{H}\big(\cdot,\mathcal{D}(t,x),\bar{U}(t,x)\big)$ is strictly concave in $U$ w.r.t. its first argument. Consequently, for a given base state $\bar{U}:[0,T]\times\Omega\to\mathbb{R}^{u}$, there exists a convex set $\mathcal{O}^{*}_{\bar{U}}\subset\mathcal{H}$ containing zero such that for every $D\in\mathcal{O}^{*}_{\bar{U}}$ one has
@@
In other words, for each fixed base state $\bar{U}:[0,T]\times\Omega\to\mathbb{R}^{u}$ there exists a convex set $\mathcal{O}^{*}_{\bar{U}}\subset\mathcal{H}$ in which the dual functional $S_{H}[D;\bar{U}]$ is convex and its Euler-Lagrange equations and side conditions are given by (3.1), with the substitution $U(t,x)\to U^{(H)}(\mathcal{D}(t,x),\bar{U}(t,x))$. The DtP map $(\mathcal{D},\bar{U})\mapsto U^{(H)}(\mathcal{D},\bar{U})$ is well-defined for $D\in\mathcal{O}^{*}_{\bar{U}}$, with the maximizer in (3.10) being an isolated critical point of $\mathcal{L}_{H}(\,\cdot\,,\mathcal{D}(t,x);\bar{U}(t,x))$.
@@
    $\mathcal{D}^{1}(t,x,\mathsf{s})\in\mathcal{O}(t,x),$ (3.15)
    $-\partial_{UU}\mathcal{L}_{H}$ $\Big(U^{(H)}(\mathcal{D}^{1}(t,x,\mathsf{s}),\bar{U}^{1}(t,x,\mathsf{s})),\mathcal{D}^{1}(t,x,\mathsf{s}),\bar{U}^{1}(t,x)\Big)>0,$
    for a.e. $(t,x)\in[0,T]\times\Omega$ and for all $\mathsf{s}\in[0,\tilde{\mathsf{s}}^{*}_{1}]$, so that, as a consequence,
@@
    or the flow reaches a fake time $\tilde{\mathsf{s}}^{*}_{1}$ when (3.15) is violated on a set of positive Lebesgue measure in $0,T]\times\Omega$.
@@
    In cases where the solutions to (3.1) are not unique, we can find only one of them. However, by varying the initial base state $\bar{U}^{1}$, it is expected to be possible to capture additional solutions as well. In principle, any solution can be retrieved, using observation (3) at the end of Sec. 3.1. In practice, see the approximation of a) a wide variety of traveling waves of a dispersive semi-discrete Burgers equation in [25] (with an algorithm which does not exploit the feature (3.14)), and b) even unstable solutions of the primal problem in a stable manner in [29, Sec. 6.3.4] that deals with the nonconvex elastodynamics of a bar (utilizing only one base state and stage).
@@
It is natural to expect that a starting base state in time-dependent problems can be more effective if the real-time interval $[0,T]$ is divided into smaller disjoint sub-intervals whose closed union covers $[0,T]$, with the dual problem solved by the above gradient flow (in fake time) in each such sub-interval, defined with the real-time initial condition generated by the solution of the previous sub-interval. Thus, in this strategy, base states are switched both in real-time (to provide constant-in-real-time starting base state guesses for each sub-interval, at a minimum), as well as potentially in fake-time for the gradient flow within each sub-interval of real time. In experience with nonlinear transient problems computed without the gradient flow strategy [23, Sec. 5.3]-[24], such switching of base states across sub-intervals in real-time was found to be essential for success (as well as computational efficiency). On the other hand, [25] involves a time-independent problem where switching base states between stages of Newton method based iterations (the analog of the gradient flow stages) was found to be essential.